# The Energy Cost of Artificial Intelligence Lifecycle in Communication Networks

### v3

@@ Abstract
Artificial Intelligence (AI) is being incorporated in several optimization, scheduling, orchestration as well as in native communication network functions. While this paradigm shift results in increased energy consumption, quantifying the end-to-end energy consumption of adding intelligence to such systems is particularly challenging. Conventional metrics focus on either communication, computation infrastructure, or model development. To address this, we propose a new metric, the Energy Cost of AI Lifecycle (eCAL) of one AI model in a system. eCAL captures the energy consumption throughout the development and deployment of an AI-model providing intelligence in a wireless communication network by analyzing the complexity of data collection and manipulation in individual components and deriving overall and per-bit energy consumption. We show that the better a model is and the more it is used, the more energy efficient an inference is. For a simple case study, eCAL for making $100$ inferences is $2.73$ times higher than for $1000$ inferences. Additionally, we have developed a modular and extendable open-source simulation tool to enable researchers, practitioners, and engineers to calculate the end-to-end energy cost with various configurations and across various systems, ensuring adaptability to diverse use cases.
Index Terms:
AI Model lifecycle, Energy Consumption, Carbon Footprint, Metric, Methodology
@@ I Introduction
Next-generation wireless networks aim to deliver unprecedented levels of connectivity, service diversity and flexibility. Artificial Intelligence (AI) is expected to become fundamental to 6G and future cellular networks [1], with traditional networking functions increasingly being supplanted by AI‑driven implementations to realize so‑called “AI‑native” capabilities  [2, 3]. These AI techniques aim to deliver intelligent, sustainable, and dynamically programmable services that enhance network adaptability and flexibility [1]. Recent works propose a pervasive multi-level native AI architecture, which incorporates knowledge graph (KG) into mobile network reducing data scale and computational costs of AI training by almost an order of magnitude [4], attempt to improve communication network efficiency by minimizing the transmitted over the air data through semantic communications [5], optimize end-to-end network slicing [6]. Furthermore, and AI/ Machine Learning (ML) workflow was defined  [7] as part of the Open Radio Access Network (O-RAN) Alliance that tackles the standardization of the network access segment.
@@
In light of climate challenges, the general AI research community’s efforts are intensifying to better estimate the costs of training [11] and using [12] AI, the Carbon Footprint (CF) of LLMs [13] and find ways to scale models efficiently [14]. Furthermore, the networking community is increasingly relying on AI techniques, are also exploring pathways towards net-zero carbon emissions [15]. They analyze the CF of different learning techniques [16] and data modalities and develop different approaches to optimize energy efficiency, including joint optimization of hardware-software co-design [17], improved scheduling [18], more efficient model design [19], and energy/carbon consumption testing [20].
Depending on the scope of the study, various metrics are employed to understand energy efficiency and CF. As well noted in [15], the traditional ITU standardized Energy-per-Bit $[J/b]$ metric ’will no longer be able to reflect the environmental impact of the modern mobile services, especially network AI-enabled smart services’. For instance, [17] considers normalized energy as well as energy for buffering in $[J]$, electricity $[Wh]$ and $[CO_{2}eq]$ for the various phases involved in federated learning (FL) edge systems [15, 16], $[J]$ and GFLOPS of the neural architectures [19] and energy $[Wh]$ for the scheduled loads in [18]. However, these metrics are not specifically designed to capture the energy efficiency of a system or its parts. To this end, a metric somewhat similar to the Energy-per-Bit $[J/b]$ would be more suitable for measuring energy efficiency of an AI-based communication system, similarly as some other efficiency metrics that have been proposed in wireless in particular and engineering and economy in general [21]. For instance, spectral efficiency, measured in $[b/s/Hz]$ [22] provides means of calculating the amount of data bandwidth available in a given amount of spectrum, and lifecycle emissions for vehicles, measured in $[g/km]$ [23], enables computing the CF in grams per kilometer.
Following these observations, we tackle a timely, unexplored and complex problem aimed to find a suitable energy efficiency metric that would be simple and general and would holistically quantify the environmental cost of inference carried out in future communication networks. This work is meant as the foundation on which to build increasingly accurate understanding of the cost of AI model embedding in networks. The contributions of this paper are as follows:
@@
  • Following the Open Systems Interconnection (OSI) and AI/ML workflow or Machine Learning Operations (MLOps) [24] architectures, we devise a methodology for determining the eCAL of an communication system and develop an open source modular and extensible simulation tool [^1] https://github.com/sensorlab/eCAL, namely, eCAL calculator. The methodology breaks down the system into different data manipulation components, and, for each component, it analyzes the complexity of data manipulation and derives the overall energy consumption.
  • Using the proposed metric, we demonstrate that the better a model is and the more it is used, the more energy-efficient an inference is. For a selected case study, the energy consumption per bit for making $100$ inferences is $2.73$ times higher than for $1000$ inferences. This can be explained by the fact that the energy cost of data collection, preprocessing, training, and evaluation, all necessary to develop an AI model are spread across more inferences.
The paper is structured as follows. We first review the relevant literature in Section II. We then describe the lifecycle of an AI model, introduce eCAL and the methodology for deriving it in Section III. Subsequently, the derivation of the energy consumption formulas for different data manipulation components involved in the AI model lifecycle are provided in Sections IV to VII. In Section VIII, the proposed metric, eCAL, is derived and analyzed for selected communication and model development cases, demonstrating its utility in capturing the overall energy cost of an AI model over its lifecycle. Finally, Section IX concludes the paper and outlines future research directions.
Table I: State-of-the-Art Energy Efficiency Metrics in Networking, Data Centers, and AI.
Category Metric Units of Measurement Description
Telecommunication Energy Efficiency $b/J$ Measures the number of information bits transmitted or received per unit of energy consumed by the radio access network or communication module. Rec. ITU-R M.2083-0 [25, 26].
Energy per Bit $J/b$ Tracks the energy consumed to transmit a single bit of information [15].
Data Center PUE % Power Usage Effectiveness (PUE) measures the ratio of the total energy used by a data center to the energy delivered to its IT equipment [27].
CUE $kg/kWh$ Carbon Usage Effectiveness (CUE) quantifies the carbon emissions (kg) produced per unit of energy consumption (kWh) in a data center [27].
WUE $L/kWh$ Water Usage Effectiveness (WUE) measures the amount of water (L) required for cooling and operations in a data center per unit of energy (kWh) supplied to IT equipment [27].
Deep Learning APC % Accuracy per Consumption (APC) measures the accuracy of a machine learning model relative to its weight energy consumption (WC). A lower WC indicates a system with reduced energy cost for the same accuracy. GreenAI-guided metric [28].
APEC % Accuracy per Energy Cost (APEC) is similar to APC but accounts for different parameters within the WC metric. GreenAI-guided metric [28].
TTCAPC % Time to Closest APC (TTCAPC) integrates training time with the APC inference metric, rewarding higher accuracies while penalizing higher net energy consumption (Wh) and longer training times. GreenAI-guided metric [28].
TTCAPEC % Time to Closest APEC (TTCAPEC) integrates training time with the APEC inference metric, rewarding higher accuracies while penalizing higher net energy consumption (measured in Euros) and longer training durations. GreenAI-guided metric [28].
Proposed Metric eCAL $J/b$ Quantifies the energy consumption of AI inference, considering both communication and computation energy costs relative to application data.
@@ II Related Work
The growing dependence on AI has introduced significant energy and carbon costs [8], intensifying environmental concerns [13]. In response, the AI research community has placed greater emphasis on understanding and then formalizing methods to assess the environmental impact of the increasing AI adoption. Due to the complexity of state-of-the-art neural architectures, in many cases the energy cost of training is computed after the training is done, by measuring the performed computation through interfaces such as the performance application programming interface [29]. Furthermore, [11] showed that the energy footprint of inference has been more extensively studied than that of training. The study highlighted that the accuracy between the predicted and measured energy consumption is below 70%.
@@
Meanwhile, a number of approaches and tools capable of proactively estimating the energy and environmental cost of training have also emerged. LLMCarbon [13] is a very recent end-to-end CF projection model designed for both dense (all parameters are used for every input) and mixture of experts (only a subset of parameters for each input) LLMs. It incorporates critical LLM, hardware, and data center parameters, such as LLM parameter count, hardware type, system power, chip area, and data center efficiency, to model both operational and embodied CFs of a LLM. Furthermore, [14] aimed to provide guidelines for scaling AI in a sustainable way. They examined the energy efficiency of new processing units and the CF of the most prominent LLM models since GPT-3, analyzed their lifecycle carbon impact, and showed that inference and training are comparable. They concluded that to enable sustainability as a computer system design principle, better tools for carbon telemetry, large-scale carbon datasets, carbon impact disclosure, and more suitable carbon metrics are required.
@@
The networking community, increasingly relying on AI techniques, are also investigating pathways towards net-zero carbon emissions [15]. In [15] the authors noticed that in spite of improvements in hardware and software energy efficiency, the overall energy consumption of mobile networks continues to rise, exacerbated by the growing use of resource-intensive AI algorithms. They introduced a novel evaluation framework to analyze the lifecycle of network AI implementations, identifying major emission sources. They proposed the dynamic energy trading and task allocation framework, designed to optimize carbon emissions by reallocating renewable energy sources and distributing tasks more efficiently across the network. Similarly to [14], they highlighted the need for the development of new metrics to quantify the environmental impact of new network services enabled by AI.
@@
The authors of [16] introduced a novel framework to quantify energy consumption and carbon emissions for vanilla FL methods and consensus-based decentralized approaches, and identified optimal operational points for sustainable FL designs. Two case studies were analyzed within 5G industry verticals: continual learning and reinforcement learning scenarios. Similar to the authors of [15], they considered energy [$Wh$] and [$CO_{2}eq$] in their evaluations. The solution proposed in [17] is a hardware/software co-design that introduces modality gating (throttling) to adaptively manage sensing and computing tasks. This so-called adaptive modality gating (AMG) solution features a novel decoupled modality sensor architecture that supports partial throttling of sensors, significantly reducing energy consumption while maintaining data flow across multimodal data: text, speech, images, and video. More energy efficient task scheduling has been considered in  [18], while more efficient neural network architecture design and subsequent model development in [19]. The energy efficiency of various programming languages was studied in [30], while energy and carbon consumption testing for AI-driven Internet of Things (IoT) services were developed in [20].
@@
Table I summarizes the relevant existing telecommunications, data center and deep learning metrics proposed to quantify the energy cost of data collection and manipulation. According to [26], standardization organizations, such as 3GPP have recently placed a greater emphasis on energy efficiency and energy savings as a core topic, and 3GPP, for instance, has listed energy efficiency as Key Performance Index (KPI) in its Release 19222https://www.3gpp.org/specifications-technologies/releases/release-19 set of specifications. While the existing energy efficiency and energy per bit metrics are seen as sufficient for previous generations of mobile communication systems, they are not able to fully capture the AI aspects estimated to also be used for implementing 6G network functionality as well as for optimizations [15]. A number of metrics such as Power Usage Effectiveness (PUE), Carbon Usage Effectiveness (CUE), and Water Usage Effectiveness (WUE) have also been proposed for data centers [27]. However, they only capture energy aspects of data center computation without providing insights into the communications and AI aspects.
@@
Furthermore, novel metrics to evaluate Deep Learning (DL) models, considering not only their accuracy and speed but also their energy consumption and cost have been proposed in [28]. The four metrics used are: Accuracy Per Consumption (APC) and Accuracy Per Energy Cost (APEC) for inference, and Time To Closest APC (TTCAPC) and Time To Closest APEC (TTCAPEC) for training. These metrics aim to promote the use of energy-efficient and green energy-powered DL systems by integrating energy considerations into the performance evaluation. Nevertheless, while powerful in terms of assessing the energy/performance trade-offs of models, these metrics are less suitable in assessing the energy costs of adding intelligence to future communication networks.
@@
Inspired by the simplicity and effectiveness of metrics such as transmission efficiency through energy-per-bit and the lifecycle emissions, and responding to calls for new metrics [15, 14] that capture the energy cost in the era of AI enabled systems, this work proposes the novel Energy Cost of AI Lifecycle (eCAL) metric and methodology to derive it. eCAL is the only metric that, based on OSI and MLOps architectures, provides insights into the cost of inference considering the data transmission and AI model overheads.
@@ III Definition and Methodology
In our work, we consider a communication and computing system that enables the development of an AI model, whose lifecycle is depicted in Fig. 1, consisting of the following data manipulating components as per the OSI stack and MLOps architectures:
@@
This data-manipulating component, depicted in the lower part of Fig. 1, includes receiving data (i.e. telemetry, applications, etc) from devices $N_{\mathrm{S}}I_{\mathrm{S}}$ to the target computing infrastructure via wired or wireless technologies. The collected information can be turned into indicators such as signal strength and be interpreted by the AI models, for example, to predict the location of a user. To successfully collect all the application-level data, the total energy of data collection $E_{\mathrm{DC}}\leavevmode\nobreak\ [J]$ is needed. The number, topology or other particularities of the data collection component of the AI model lifecycle can be adapted to specific configurations if needed.
@@
To ensure accuracy and reliability during the training process, the data must go through several preprocessing steps, such as cleaning, feature engineering and transformation. The energy consumption for preprocessing, $E_{\mathrm{pre}}$, depends on the integrity of the ingested dataset. The metric is conceptually capable of support also distributed or hybrid pre-processing as well as training and evaluation scenarios.
@@
The training component comprises the AI/ML model development using selected AI/ML techniques, such as neural network architectures, and the training data. In this step of the model development process, the processed data, $N_{\mathrm{S,T}}$ is fetched and utilized to learn weights and biases in order to approximate the underlying distribution. For most neural architectures, the learning processes rely on Graphics Processing Units and Tensor Processing Units, which perform complex tensor processing operations and consuming training energy, $E_{\mathrm{train}}$. Once the neural network architecture weights are learned using the data in view of minimizing a loss function, the model is considered ready for evaluation and deployment. Subsequently, the quality of the learned model is evaluated on the evaluation dataset, $N_{\mathrm{S,E}}$, a process that consumes $E_{\mathrm{eval}}$ energy. The data storage, preprocessing, training, and evaluation form the end-to-end training of the AI/ML model and cumulatively require $E_{\mathrm{D}}$ energy to complete.
@@
Once the model is trained, it can be used by applications or within network scheduling and optimization components in inference mode. Samples of data can be sent to the AI model running in inference mode $N_{\mathrm{I,P}}$ and model outputs are returned as responses in the form of forecasts for regression problems or discrete (categorical) labels for classification problems. The energy consumption $E_{\mathrm{inf}}$ of a single inference is relatively small, however, for high volumes of requests, it can become significant.
@@
$eCAL$ EC per bit over the lifetime of an AI model in the system
@@
Based on this lifecycle, we propose eCAL as the ratio of total energy consumed by data manipulation components to the total manipulated application-level bits:
$eCAL$ $=$
Total energy of data manipulation components [$\frac{\text{Total energy of data manipulation components }[J]}{\text{Total manipulated application level data }[b]}.$ (1)
In the upcoming sections, we employ the following methodology in order to derive the proposed metric for example case studies that comply with the OSI and MLOps architectures. We follow the data flow, highlighted with blue in Fig. 1, across the above-mentioned components and derive the computational complexity and subsequently the energy consumption, highlighted with purple, in each component. For instance, in Section IV, we derive $E_{\mathrm{T}}$ for the data collection component. With this formalism in place, Section VIII is concerned with deriving eCAL under specific configuration assumptions, without limiting the scope of the proposed metric.
For the sake of clarity and convenience, we summarize the notations used throughout this paper in Table II.

### v4

@@ Abstract
Artificial Intelligence (AI) is being incorporated in several optimization, scheduling, orchestration as well as in native communication network functions. This paradigm shift results in increased energy consumption, however, quantifying the end-to-end energy consumption of adding intelligence to communication systems remains an open challenge since conventional energy consumption metrics focus on either communication, computation infrastructure, or model development. To address this, we propose a new metric, the Energy Cost of AI Lifecycle (eCAL) of an AI model in a system. eCAL captures the energy consumption throughout the development, deployment and utilization of an AI-model providing intelligence in a communication network by (i) analyzing the complexity of data collection and manipulation in individual components and (ii) deriving overall and per-bit energy consumption. We show that as a trained AI model is used more frequently for inference, its energy cost per inference decreases, since the fixed training energy is amortized over a growing number of inferences. For a simple case study we show that eCAL for $100$ inferences is $2.73$ times higher than for $1000$ inferences. Additionally, we have developed a modular and extendable open-source simulation tool to enable researchers, practitioners, and engineers to calculate the end-to-end energy cost with various configurations and across various systems, ensuring adaptability to diverse use cases.
@@
GSMA
Global System for Mobile communications Alliance
RAN
radio access network
@@
NIC
Network Interface Card
DMA
Direct Memory Access
PC
Personal Computer
@@ I Introduction
Next-generation wireless networks aim to deliver unprecedented levels of connectivity, service diversity and flexibility. Artificial Intelligence (AI) is expected to become fundamental to 6G and future cellular networks [1], with traditional networking functions increasingly being supplanted by AI‑driven implementations to realize so‑called “AI‑native” capabilities [2, 3]. These AI techniques aim to deliver intelligent, sustainable, and dynamically programmable services that enhance network adaptability and flexibility [1]. Recent works propose a pervasive multi-level native AI architecture that incorporates knowledge graph (KG) into mobile network reducing data scale and computational costs of AI training by almost an order of magnitude [4], attempt to improve communication network efficiency by minimizing the transmitted data through semantic communications [5], and optimize end-to-end network slicing [6]. Furthermore, and AI/ Machine Learning (ML) workflow was defined [7] as part of the Open Radio Access Network (O-RAN) Alliance that tackles the standardization of the network access segment.
@@
In light of climate challenges, the general AI research community’s efforts are intensifying. They aim to better estimate the costs of training [11] and using AI [12], assess the Carbon Footprint (CF) of LLMs [13], and find ways to scale models efficiently [14]. Furthermore, while increasingly relying on AI techniques, the networking community is also exploring pathways towards net-zero carbon emissions [15]. In particular, researchers analyze the CF of different learning techniques [16] and data modalities and develop different approaches to optimize energy efficiency, including joint optimization of hardware-software co-design [17], improved scheduling [18], more efficient model design [19], and energy/carbon consumption testing [20].
Depending on the scope of the study, various metrics are employed to understand energy efficiency (EE) and CF. As well noted in [15], the traditional ITU standardized Energy-per-Bit $[J/b]$ metric ’will no longer be able to reflect the environmental impact of the modern mobile services, especially network AI-enabled smart services’. For instance, [17] considers both normalized energy and energy for buffering (in $[J]$), whereas [15, 16] examine electricity consumption (in $[Wh]$) and carbon emissions (in $CO_{2}eq]$) for the various phases involved in federated learning (FL) edge systems. The energy cost and computational complexity ($[J]$ and GFLOPS) of neural architectures are evaluated in [19], while [18] focuses on energy cost $[Wh]$ of scheduled loads. However, these metrics are not specifically designed to capture the EE of a system or its parts. To this end, a metric somewhat similar to the Energy-per-Bit $[J/b]$ would be more suitable for measuring EE of an AI-based communication system, similarly as some other efficiency metrics that have been proposed in wireless systems in particular and in engineering and economy in general [21]. For instance, spectral efficiency, measured in $[b/s/Hz]$ [22] provides means of calculating the amount of data bandwidth available in a given amount of spectrum, and lifecycle emissions for vehicles, measured in $[g/km]$ [23], enable computing the CF in grams per kilometer.
Following these observations, we tackle a timely, unexplored and complex problem aimed to find a suitable EE metric that would be simple and general and would holistically quantify the environmental cost of inference carried out in future communication networks. This work is meant as the foundation on which to build increasingly accurate understanding of the cost of AI model embedding in networks. The contributions of this paper are as follows:
@@
  • Following the standard Open Systems Interconnection (OSI) model and AI/ML workflow or Machine Learning Operations (MLOps) [24], falling under the platform layer of the standard Cloud Computing Reference Architecture (CCRA), we devise a methodology for determining the eCAL of an communication system and develop an open source modular and extensible eCAL calculator [^1] https://github.com/sensorlab/eCAL. The methodology breaks down the system into different data manipulation components, and, for each component, it analyzes the complexity of data manipulation and derives the overall energy consumption.
  • Using the proposed metric, we demonstrate that the better a model and the more it is used, the more energy-efficient is inference. For a selected case study, the energy consumption per bit for $100$ inferences is $2.73$ times higher than for $1000$ inferences. This can be explained by the fact that the energy cost of data collection, preprocessing, training, and evaluation, all necessary to develop an AI model, are spread across more inferences.
The paper is organized as follows. Section II reviews related work. Section III describes the AI model lifecycle, introduces eCAL, and outlines its derivation methodology. Sections IVVII present energy consumption formulas for the data manipulation components of the lifecycle. Section VIII derives and analyzes eCAL for selected communication and model development cases, illustrating its effectiveness in capturing overall lifecycle energy costs. Section IX concludes the paper and discusses future research directions.
@@ II Related Work
The growing dependence on AI has introduced significant energy and carbon costs [8], intensifying environmental concerns [13]. In response, the AI research community has placed greater emphasis on understanding and then formalizing methods to assess the environmental impact of the increasing AI adoption. Due to the complexity of state-of-the-art neural network architectures, in many cases the energy cost of training is computed after the training is done, by measuring the performed computation through interfaces such as the performance application programming interface [25]. Furthermore, [11] showed that the energy footprint of inference has been more extensively studied than that of training. The study highlighted that the accuracy between the predicted and measured energy consumption is below 70%.
Emerging in the AI Community
@@
A number of approaches and tools capable of proactively estimating the energy and environmental cost of training have also emerged. LLMCarbon [13] is a very recent end-to-end CF projection model designed for both dense (all parameters are used for every input) and mixture of experts (only a subset of parameters for each input) LLMs. It incorporates critical LLM, hardware, and data center parameters, such as LLM parameter count, hardware type, system power, chip area, and data center efficiency, to model both operational and embodied CFs of an LLM. Furthermore, [14] aimed to provide guidelines for scaling AI in a sustainable way. They examined the EE of new processing units and the CF of the most prominent LLM models since GPT-3, analyzed their lifecycle carbon impact, and showed that inference and training are comparable. They concluded that to enable sustainability as a computer system design principle, better tools for carbon telemetry, large-scale carbon datasets, carbon impact disclosure, and more suitable carbon metrics are required.
Table I: State-of-the-art energy efficiency metrics, their relevance across the AI lifecycle, and additional system context.
Column (2) (3) (4) (5) (6) (7) (8) (9) (10) (11)
Characteristic Metrics
Energy Efficiency Energy per Bit PUE CUE WUE APC APEC TTCAPC TTCAPEC eCAL
Category Telecom Telecom Data Center Data Center Data Center Deep Learning Deep Learning Deep Learning Deep Learning Proposed
Units $b/J$ $J/b$ % $kg/kWh$ $L/kWh$ % % % % $J/b$
Description Measures information bits per Joule consumed by the radio access network or communication module. Energy consumed to transmit one bit over the link. Power Usage Effectiveness: ratio of total facility energy to energy delivered to IT. Carbon Usage Effectiveness: CO2 ($kg$) per kWh consumed. Water Usage Effectiveness: liters of water per kWh for cooling/operation. Accuracy per Consumption: accuracy relative to the energy cost of model weights (GreenAI). Accuracy per Energy Cost: variant of APC capturing broader costs. Time to Closest APC: integrates training time and energy into APC. Time to Closest APEC: integrates training time and energy into APEC. End-to-end energy of AI data collection/ inference relative to application bits (comms + compute).
Mapping to the telco OSI standard layers and cloud computing CCRA standard layers.
OSI Standard Physical Physical All All All Application Application Application Application All
CCRA Standard All All All Service Service Service Service Service, Platform
Relevance to the data-manipulating components in the AI lifecycle
Data Collection (Sec. IV)
Preproc. (Sec. V)
Training (Sec. VI)
Evaluation (Sec. VII)
Inference (Sec. VIII)
Emerging in the Networking Community
@@
Increasingly relying on AI techniques, the networking community is also investigating pathways towards net-zero carbon emissions [15]. In [15] the authors noticed that in spite of improvements in hardware and software energy efficiency, the overall energy consumption of mobile networks continues to rise, exacerbated by the growing use of resource-intensive AI algorithms. They introduced a novel evaluation framework to analyze the lifecycle of network AI implementations, identifying major carbon emission sources. They proposed the dynamic energy trading and task allocation framework, designed to optimize carbon emissions by reallocating renewable energy sources and distributing tasks more efficiently across the network. Similar to [14], they highlighted the need for the development of new metrics to quantify the environmental impact of new network services enabled by AI.
@@
The authors of [16] introduced a novel framework to quantify energy consumption and carbon emissions for vanilla FL methods and consensus-based decentralized approaches, and identified optimal operational points for sustainable FL designs. Two case studies were analyzed within 5G industry verticals: continual learning and reinforcement learning scenarios. Similar to the authors of [15], they considered energy [$Wh$] and carbon emissions $CO_{2}eq$] in their evaluations. The solution proposed in [17] is a hardware/software co-design that introduces modality gating (throttling) to adaptively manage sensing and computing tasks. It features a novel decoupled modality sensor architecture that supports partial throttling of sensors, significantly reducing energy consumption while maintaining data flow across multimodal data: text, speech, images, and video. More energy efficient task scheduling has been considered in [18], while more efficient neural network architecture design and subsequent model development in [19]. EE of programming languages was studied in [26], while energy and carbon consumption testing for AI-driven Internet of Things (IoT) services were developed in [20].
Existing Metrics and Scope Mapping to OSI and CCRA standards
@@
Table I summarizes the relevant existing telecommunications (columns 2 and 3), data center (columns 4-6), and deep learning (columns 7-10) metrics proposed to quantify the energy cost of data collection and manipulation. The telecommunication metrics are well established, however they measure EE and energy per bit respectively on the physical layer as summarized in the table. These EE metrics were sufficient for previous generations of mobile communication systems, but they cannot capture the AI aspects to be used for implementing 6G network functionality as well as for optimizations [15]. Furthermore, according to [27], standardization organizations have recently placed a greater emphasis on EE and energy savings as a core topic, and 3GPP, for instance, has listed EE as Key Performance Index (KPI) in its Release 19222https://www.3gpp.org/specifications-technologies/releases/release-19 set of specifications.
The data center consumption metrics in Table I, i.e. Power Usage Effectiveness (PUE), Carbon Usage Effectiveness (CUE), and Water Usage Effectiveness (WUE) [28], refer to the overall data center operation and include computing for AI and non AI purposes, networking and cooling. They offer a macro-system view agnostic to specifics of AI or communications as can be seen from their description in the table (see row ”description”). They implicitly quantify the energy consumption across all layers of the OSI and CCRA as shown in the rows labeled OSI Standard and CCRA Standard, but they are not designed to provide per layer analytical insights and the energy per amount of processed information.
@@
The Deep Learning (DL) metrics summarized in columns 7-10, namely Accuracy Per Consumption (APC), Accuracy Per Energy Cost (APEC) for inference, Time To Closest APC (TTCAPC) and Time To Closest APEC (TTCAPEC) [29], consider not only the accuracy and speed of DL models but also their energy consumption and cost for training. Nevertheless, while powerful in terms of assessing the energy/performance trade-offs of models, these metrics are less suitable in assessing the energy costs of adding intelligence to future communication networks, especially in cases when the dominant energy consumption arises from data collection rather than from the model training/inference. As marked in the table (OSI Standard and CCRA Standard), their scope falls under the application layer in OSI and service layer in CCRA.
Summary
@@
Inspired by the simplicity and effectiveness of metrics such as transmission efficiency through energy-per-bit and the lifecycle emissions, and responding to calls for new metrics [15, 14] that capture the energy cost in the era of AI enabled systems, this work proposes the novel Energy Cost of AI Lifecycle (eCAL) metric and methodology to derive it. eCAL is the only metric that, based on OSI and CCRA, provides insights into the cost of inference, considering the data transmission and AI model overheads.
@@ III Definition and Methodology
In our work, we consider a communication and computing system that enables the development of an AI model, whose lifecycle is depicted in Fig. 1, consisting of the following data manipulating components as per OSI and MLOps:
@@
This data-manipulating component, depicted in the lower part of Fig. 1, represents receiving $N_{\mathrm{S}}$ data samples with $I_{\mathrm{S}}$ sample size from distributed devices via $N_{\mathrm{L}}$ independent wired or wireless links at the target computing infrastructure. The collected information can be turned into indicators such as signal strength and be interpreted by the AI models, for example, to predict the location of a user. To successfully collect all the application-level data, the total energy of data collection $E_{\mathrm{DC}}\penalty 10000\ [J]$ is needed. The number, topology or other particularities of the data collection component of the AI model lifecycle can be adapted to specific configurations if needed.
@@
To ensure accuracy and reliability during the training process, the data must go through several preprocessing steps, such as cleaning, feature engineering and transformation. The energy consumption for preprocessing, $E_{\mathrm{pre}}$, depends on the integrity of the ingested dataset. The proposed metric is conceptually capable of supporting also distributed or hybrid pre-processing as well as training and evaluation scenarios.
@@
The training component comprises the AI/ML model development using selected AI/ML techniques, such as neural network architectures, and the training data. In this step of the model development process, the processed data, $N_{\mathrm{S,T}}$ is fetched and utilized to learn weights and biases in order to approximate the underlying distribution. For most neural network architectures, the learning processes rely on Graphics Processing Units and Tensor Processing Units, which perform complex tensor processing operations and consume training energy, $E_{\mathrm{train}}$. Once the neural network architecture weights are learned using the data in view of minimizing a loss function, the model is considered ready for evaluation and deployment. Subsequently, the quality of the learned model is evaluated on the evaluation dataset, $N_{\mathrm{S,E}}$, in a process that consumes $E_{\mathrm{eval}}$ energy. The data storage, preprocessing, training, and evaluation form the end-to-end training of the AI/ML model and cumulatively require $E_{\mathrm{D}}$ energy to complete.
@@
Once the model is trained and evaluated, it can be used by applications or within network scheduling and optimization components in inference mode. $N_{\mathrm{I,P}}$ samples of data can be sent to the AI model running in inference mode and model outputs are returned as responses in the form of forecasts for regression problems, or discrete (categorical) labels for classification problems. The energy consumption $E_{\mathrm{inf}}$ of a single inference is relatively small; however, for high volumes of requests, it can become significant.
@@
eCALeCAL EC per bit over the lifetime of an AI model in the system
@@
III-A Existing Metric Relevance for the Data Manipulating Components of the AI Lifecycle
Existing energy metrics offer only partial visibility into the energy demands of AI models across their full lifecycle, as illustrated in Table I. On the one hand, telecommunication and data center metrics (columns 2–6) primarily address the data collection stage. Telecommunication metrics focus on physical-layer energy consumption in the access network, while data center metrics capture aggregate energy use from all IT infrastructure. However, these metrics lack formal attribution to specific network segments or OSI layers, limiting their granularity and diagnostic utility. On the other hand, the next four columns (columns 7–10) show that existing deep learning metrics provide only coarse coverage of energy consumption during data preprocessing, training, evaluation, and inference. These metrics are typically limited to classification tasks and often rely on weight counts as proxies for energy use. Importantly, only two of them explicitly consider inference, underscoring a significant gap in coverage. Therefore, there is a need for the development and adoption of the proposed eCAL metric (column 11) as a unified, lifecycle-aware measure that enables consistent and component-resolved evaluation of energy efficiency in AI systems.
III-B Proposed Metric
Based on the AI model lifecycle illustrated in Figure 1 and discussed in this section, we propose eCAL as the ratio of total energy consumed by data manipulation components to the total manipulated application-level bits:
[…]
eCAL introduces a fundamentally new way of looking and assessing the cost of using AI - it essentially enables to quantify the energy cost of each bit that comes out from the inference of an AI model over its entire lifetime while accounting for the underlying overheads. The purpose of eCAL is to measure the overall energy consumption of any AI model developed using data and a neural network architecture, independent of the use case and accuracy/performance expected by that use case, similarly as for instance, EE from Table I is applicable to various communication modules independent on their spectral efficiency. As illustrated in Table I, eCAL is the only metric able to capture the energy costs in communication networks following the standard OSI model by modeling the overheads introduced by various layers as well as two of the layers of the standard CCRA in cloud computing. All the other metrics have a narrower scope and do not center on and include the actual AI training and inference. In future when a metric such as eCAL may be standardized, it could be used as an energy badge on any AI model, in a similar fashion as there are CO2 footprint badges used when booking flights or EE classes used on household appliances.
III-C Methodology
In the upcoming sections, we derive the proposed metric for example case studies that comply with the OSI standard architecture and MLOps framework that is part of the platform layer in the CCRA standard architecture. We follow the data flow, highlighted with blue in Fig. 1, across the above-mentioned components and derive the computational complexity and subsequently the energy consumption, highlighted with purple, in each component. In Section IV, we derive $E_{\mathrm{T}}$ for the data collection component. With this formalism in place, Section VIII is concerned with deriving eCAL under specific configuration assumptions, without limiting the scope of the proposed metric. For the sake of clarity and convenience, we summarize the notations used throughout this paper in Table II.